The ratio of the interior angle to the exterior angle of a polygon is 5:2.how many sides has the polygon
Answer by Alan3354(69443)

The ratio of the interior angle to the exterior angle of a polygon is 5:2.how many sides has the polygon
--------------
5x + 2x = 180
x = 180/7
----------
Ext angles = 2x = 360/7
# of sides = 360/Ext
# of sides = 7